OpenStudy (anonymous):

PDE's: Solve the following PDE using separation of variables: 3u_x + 2u_y = 0, u(x,0) = 4e^(-x).

OpenStudy (anonymous):

Please confirm, I got u(x,y) = 4e^(-x+y)

OpenStudy (anonymous):

I am not getting a zero when I plug in your answer

OpenStudy (anonymous):

you can check the answer by taking derivatives and seeing if it satifies the equation. Ux = -4e^(y-x) Uy=4e^(y-x) 3*-4e^(y-x)+2*4e^(y-x) = -4e(y-x), so that can't be right. The answer is 4e^(1.5y-x)

OpenStudy (anonymous):

u(x,y)=X(x)Y(y) u_x= X' Y u_y= X Y' 3X' Y=-2 X Y' \[\frac{-3}{2} \frac{X'}{X}=\frac{Y'}{Y}\] \[\frac{-3}{2} \frac{X'}{X}=C\] \[-3/2 log X =C x\] \[X = e^{\frac{-(2 C x)}{3}}\] \[\frac{Y'}{Y}=C\] \[Log[Y]=Cy+D\] \[Y= K e^{c y}\]

OpenStudy (anonymous):

u(x,y)= \[Ke^{cy} e^{(-2/3) C x}\] using condition K=4 C= -3/2 u(x,y)= \[4 e^{3/2y} e^{ x}\]

OpenStudy (anonymous):

sorry C= 3/2

OpenStudy (anonymous):

oh okay i see what I did wrong. I made both coefficients of x and y as 1 instead of just one of them being 1 and the other 3/2. thanks.

OpenStudy (anonymous):

(careless mistake sorry, lol)

OpenStudy (anonymous):

are you taking PDE course?

OpenStudy (anonymous):

almost. the course is 3/4 linear algebra and the last 1/4 is PDE's and Fourier Series

OpenStudy (anonymous):

Fourier Series are the best

OpenStudy (anonymous):

lol haha I could have taken the other alternative, which has the last 1/4 as more on vector calculus (Green's, stoke's and stuff). what would have been better for a computer engineering course? my current one or that other alternative?

OpenStudy (anonymous):

I'm sorry I couldn't quite understand that. are not as differential, and linear algebra?

OpenStudy (anonymous):

In my opinion as, Electrical Engineering student, (Green's, stoke's and stuff) are not as useful as differential, and linear algebra

OpenStudy (anonymous):

oh alright, sorry. but yeah, I've seen terms like Fast Fourier Transform and signal optimization in my introductory engineering courses, so I kind of have an idea how the last 1/4 could apply to my major. linear algebra has some nice computer applications too like in search engines and stuff, i think? haha.

OpenStudy (anonymous):

Yeah, but that's mainly Electrical Engineering material , not sure if Computer Engineering students have to take Signal processing?

OpenStudy (anonymous):

the design of my major is closer to the side of electrical engineering than computer science in the spectrum, so yeah, I do have to take a course on signals and systems.

OpenStudy (anonymous):

Signal processing is math intensive Fourier, Laplace, ...

OpenStudy (anonymous):

oh wow.. haha better ready myself for my junior year then

OpenStudy (anonymous):

Taken physics already?

OpenStudy (anonymous):

yeah I've already finished the freshman level general subjects in my community college

OpenStudy (anonymous):

oh good, because it's killing me

OpenStudy (anonymous):

haha which one are you in?

OpenStudy (anonymous):

E&M

OpenStudy (anonymous):

oh yeah that one almost killed me too. you can do it, it's easier than dealing with PDE's and fourier, at least for me lol

OpenStudy (anonymous):

Where do you go to school?

OpenStudy (anonymous):

Texas A&M